OpenStudy (anonymous):

what is 1 over 14 rounded to the hundredths place

OpenStudy (anonymous):

Long division time!

OpenStudy (anonymous):

long division to 3 decimal places, then round

OpenStudy (anonymous):

or you could cheat and find a calculator

OpenStudy (anonymous):

In my classroom, were learning how to convert fractions to decimals and one of the questions was how do you round 1over 14 and I didn't know if the number would stay the same or long division it.

OpenStudy (anonymous):

long division or use a calculator would give you the same answer

OpenStudy (anonymous):

personally you should use long division for the practice

OpenStudy (anonymous):

+1 for "you should use long division for the practice"

OpenStudy (anonymous):

what is 0.07 as a percent?

OpenStudy (anonymous):

what is 0.07*100

OpenStudy (anonymous):

to be honest I really don't know. :(

OpenStudy (anonymous):

ok then 1*100=100 .1*100=10 .01*100=1 .001*100=.1 .0001*100=.01 notice the decimal place

OpenStudy (anonymous):

So the answer to the other question would be .0007*100=.07?

OpenStudy (anonymous):

well you have .07*100, so that would just be 7

OpenStudy (anonymous):

so it would just be 7% I'm taking a guess here.